a quadratic graph has minimum point (-1,2). find the equation of the graph
Answer by nerdybill(7384)

Reference:
http://www.mathwarehouse.com/geometry/parabola/standard-and-vertex-form.php
.
The vertex form of a parabola:
y= a(x-h)^2+k
.
Where
(h,k) is the vertex
.
Here they give you the vertex as:
(h,k) = (-1,2)
.
y= a(x-(-1))^2+2
And since the problem stated that is was a MINIMUM... 'a' is POSITIVE.
y= (x+1)^2+2
y= x^2+2x+1+2
y= x^2+2x+3